# Risk-sensitive Reinforcement Learning Based on Convex Scoring Functions

### v1

@@ Abstract
Keywords: Risk-sensitive reinforcement learning, Actor–Critic algorithm, convex scoring function, augmented state process, two-stage optimization, error analysis
@@ 1 Introduction
However, due to the time non-separable nature in (2), it is well known that the time-inconsistency issue arises so that the standard dynamic programming principle cannot be employed directly in devising the algorithm. Our technical and practical contributions are summarized below.
@@
Secondly, we develop an Actor-Critic (AC) algorithm to optimize the policy and value function using neural networks. The Critic-step ensures the value function approximates the true value uniformly, while the Actor-step updates the policy based on the policy gradient. Theoretical results demonstrate the approximation of the algorithm and provide bounds on the approximation error for both the value function (Theorem 2) and the optimal auxiliary variable $\upsilon$ (Theorem 3).
@@ 2.2 Risk Measures and Scoring Functions
In the context of quantitative risk management, a scoring function is a metric to evaluate and compare competing risk forecasts. As a definition (Chapter 9 of McNeil et al., (2015)), a classical scoring function $f:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ satisfies
  • $f(y,\upsilon)\geqslant 0$ and $f(y,\upsilon)=0$ if and only if $y=\upsilon;$
  • $f(y,\upsilon)$ is increasing in $\upsilon$ for $\upsilon>y$ and decreasing for $\upsilon<y;$
  • $f(y,\upsilon)$ is continuous in $\upsilon.$
@@
Definition 1 is quite general because the convexity is closely tied to unimodality. If the convexity is replaced by the condition that $-f(y,\upsilon)$ is unimodal in $\upsilon$ (i.e., for fixed $y$, there exists some $\upsilon_{0}\in\mathbb{R}$ such that $f(y,\cdot)$ is decreasing on $(-\infty,\upsilon_{0}]$ and increasing on $[\upsilon_{0},\infty)$), then the scoring function satisfies the second condition in the definition of the classical scoring function above; see also Bellini and Bignozzi, (2015). In our paper, we focus on the convexity-based formulation to enhance the clarity and facilitate some mathematical arguments. The next result for stochastic optimization readily follows from Definition 1.
@@ 3 Problem Formulation
where the function $f$ is a convex scoring function and $h$ is a bivariate transform function strictly increasing w.r.t. its first argument, and the second argument $\upsilon$ is the auxiliary variable. Such risk measures are not required to be coherent or even monetary, but they encompass a broad class of functionals and many cases of coherent risk measures, such as expectation $\rho(Y)=\min_{\upsilon\in\mathbb{R}}\mathbb{E}[Y]$ and ES $\rho(Y)=\min_{\upsilon\in\mathbb{R}}\mathbb{E}[\upsilon+\frac{1}{\alpha}(Y-\upsilon)^{+}].$ Many convex (but not coherent) risk measures are also covered, such as entropic risk measures $\rho(Y)=\frac{1}{\gamma}\log\left(\mathbb{E}[\exp(\gamma Y)]\right)$ with a risk-aversion parameter $\gamma>0$ and expected utility $\rho(Y)=\mathbb{E}[u(Y)]$ with a general convex function $u$ (see Rockafellar and Uryasev, (2013)). Note that some risk measures in our consideration are not even monetary, such as median absolute deviation (MAD) $\rho(Y)=\min_{\upsilon\in\mathbb{R}}\mathbb{E}[|Y-\upsilon|],$ variance $\rho(Y)=\min_{\upsilon\in\mathbb{R}}\mathbb{E}[|Y-\upsilon|^{2}]$ and asymmetric variance $\rho(Y)=\inf_{\upsilon\in\mathbb{R}}\,\mathbb{E}\left[\alpha(Y-\upsilon)^{2}\cdot\mathds{1}_{\left\{Y\geqslant\upsilon\right\}}+(1-\alpha)(Y-\upsilon)^{2}\cdot\mathds{1}_{\left\{Y\leqslant\upsilon\right\}}\right]$. Another important class of risk measures is the family of mean-risk utilities, which are constructed as the combination of an expectation and a risk penalty.
@@ 4 Methodology
We need to further restrict the range of $\upsilon$ to a compact set such that it contains at least one optimal solution $\upsilon^{*}$ of Problem (7). To this end, we first show in Proposition 1 below that the range of $\upsilon$ can be restricted to a compact set. However, in some cases, the conditions in Proposition 1 may not be satisfied (see Remark 3) and we have to assume that the minimum of the objective is attained on $[\underline{\upsilon},\bar{\upsilon}]$ for some pre-known $\underline{\upsilon},\bar{\upsilon}\in\mathbb{R}.$ This assumption is commonly used in some existing studies in reinforcement learning (see Baxter and Bartlett, (2001)). Note that even when this assumption is not fulfilled, we can still find $\varepsilon$-correct range for $\upsilon,$ i.e., that for any $\varepsilon>0,$ there exists an interval $[\underline{\upsilon}^{\varepsilon},\bar{\upsilon}^{\varepsilon}]$ such that
@@
for any random variable $Y$ satisfying $\underline{y}\leqslant Y\leqslant\bar{Y}$ almost surely. To address a similar issue, Ni and Lai, (2022) derives a lower bound of the optimal $\upsilon$ (which agrees with Remark 3) and imposes an assumption on its upper bound for the entropic VaR risk measure.
@@
Consider the entropic VaR mentioned in Example 2. With $f(y,\upsilon)=\frac{1}{\alpha}\exp(e^{\upsilon}y)$, $h(x,\upsilon)=e^{-\upsilon}\log(x)$ and $g(x)=\exp(x),$ condition holds for the entropic VaR if $[\underline{y},\bar{y}]\subset[0,\infty)$. However, conditions (2) and (3) are not satisfied because $g(h(f(y,\upsilon),\upsilon))$ is strictly decreasing for $\upsilon\in\mathbb{R}$ for fixed $y.$ In this case, for any $\varepsilon>0,$ we can find $[\underline{\upsilon}^{\varepsilon},\bar{\upsilon}^{\varepsilon}]$ such that
@@
The idea behind this formula is similar to the approach proposed by Bäuerle and Ott, (2011) to reduce the MDP under ES. However, their approach relies on the specific form of $(y-\upsilon)^{+}$ in the scoring function for the ES risk measure, which facilitates the simple fusion of $Y$ and $\upsilon$ into one state process. For the convex scoring function, their approach no longer holds. However, our proposed method separates the variable $\upsilon$ and the augmented state process $Y$ such that we can devise the RL algorithm to accommodate more types of risk measures.
@@
Combining Theorem 1, the dominated convergence theorem, and the continuity of $h,$ the next proposition holds.
@@
Therefore, studying the properties of $\mathbb{E}[V^{*}_{0,\upsilon}(S_{0},0)]$ is of fundamental importance. Proposition 3 establishes the continuity of $\upsilon\mapsto\mathbb{E}[V^{*}_{0,\upsilon}(S_{0},0)]$, a key result in showing the approximation of our RL algorithm in later subsections. Notably, our analysis does not require any assumption of continuity on the MDP, rendering the proof of Proposition 3 technically nonstandard.
@@
$\upsilon\mapsto\mathbb{E}[V^{*}_{0,\upsilon}(S_{0},0)]$ is continuous on $\Upsilon$ and $\upsilon\mapsto h(\mathbb{E}[V^{*}_{0,\upsilon}(S_{0},0)],\upsilon)$ has a minimizer (not necessarily unique) on $\mathbb{R}.$ Furthermore,
@@ 4.1 Part 1: Gradient Calculation and Neural Network (NN) Training
We adopt the policy gradient Actor-Critic RL algorithm for the inner optimization problem; see Konda and Tsitsiklis, (2003). That is, for the inner dynamic optimization problem, we alternatively learn two functions, the value function $V$ and the policy $\pi.$ In our work, both the value function and the policy are characterized by fully connected and multilayered feed-forward networks, for which the parameters are denoted by $\theta\in\Theta$ and $\phi\in\Phi$, respectively. The inputs to both networks include $(t,\upsilon,s,y).$ The network’s outputs are continuous with respect to $(t,\upsilon,s,y).$ See also some existing studies using the policy gradient approach on SRP in Tamar et al., (2016), Zhang et al., (2021), Ni and Lai, (2022)) and Coache and Jaimungal, (2024)). In contrast to these studies, our algorithm does not require the simulation-upon-simulation approach or the sample-average-approximation approach, which greatly reduces the computational complexity in simulating trajectories.
@@ 4.3 Part 3: Sampling the auxiliary variable $\upsilon$
Perform L(n) critic steps and get $\text{Perform $L(n)$ critic steps and get $\phi_{n}$};$
Perform L(n) actor steps and get $\text{Perform $L(n)$ actor steps and get $\theta_{n}$};$
@@
we have $\lim_{n\to\infty}h(\mathbb{E}[V^{\pi_{n,\varepsilon}}_{0,\upsilon_{n+1}}(S_{0},0)],\upsilon_{n+1})$ exists with a finite value. Furthermore, if there exists a nonnegative function $\delta:\Upsilon\times\Upsilon\to[0,\infty)$ such that for any $\upsilon^{\prime}\in\Upsilon,\ \pi\in\tilde{\Pi},$ the following two conditions hold:
@@
Algorithm 4 The Complete Recipe based on Convex Scoring Function
@@
4.3.1 Expected Shortfall Case
Expected shortfall ($\mathrm{ES}$) is widely used in risk-sensitive RL (see Bäuerle and Ott, (2011), Chow et al., (2015), Tamar et al., (2015), Kashima, (2007), Du et al., (2022), Godbout et al., (2021)). Our approach can handle the $\mathrm{ES}$ objective by taking $f(y,\upsilon)=\upsilon+(y-\upsilon)^{+}$ and $h(x,\upsilon)=x.$ We present here that under the $\mathrm{ES}$ objective, the sampling method described above can be further simplified to avoid the iterative use of SGD to solve (14). For any initial $\upsilon^{*}_{0}$ and parameters $\phi_{0}$ and $\theta_{0},$ we sample $\upsilon$ and update the parameters according to the following process: in stage $n\ (n\geqslant 1),$ we
@@ 4.3.1 Expected Shortfall Case
Perform L(n) critic steps and get $\text{Perform $L(n)$ critic steps and get $\phi_{n}$};$
Perform L(n) actor steps and get $\text{Perform $L(n)$ actor steps and get $\theta_{n}$};$
@@
It holds that $\lim_{n\to\infty}\mathbb{E}[V^{\pi_{n,\varepsilon}}_{0,\upsilon_{n+1}}(S_{0},0)]$ exists with a finite value. Furthermore, if there exists some $\Delta>0$ such that for any $\upsilon^{\prime}\in\Upsilon,\ \pi\in\tilde{\Pi},$ the next two conditions hold:
@@ 5 Application: Statistical Arbitrage
  • RL-$\mathrm{Mean}$-$\mathrm{Var}:$ our proposed RL model minimizing the mean-variance utility, i.e. $\rho(Y)=\mathbb{E}[Y]+\lambda\mathbb{E}[(Y-\mathbb{E}Y)^{2}],$ where $\lambda$ is chosen as $1$;
@@ 5.2 Policy Interpretation (Initial and Dynamic)
Figure 4 shows the policy learned after time $0$ Since the proposed method treats $y$ as a new (augmented) state, to understand the learned strategy, we propose a three-dimensional heat map illustration, which represents the action on different $(y,S,q)$ triples. Similar to time $0,$ the agent still suggests buying when the price is low and selling when the price is high again, but the policy becomes conservative as the value of $y$ increases. This can be illustrated with an example at $t=2$ in Figure 4, noting that when the price is very low (close to 0.5), with the same inventory $Q=2,$ the agent buys when the $y$ value is smaller and tends to maintain the position when the $y$ value is larger. Another point is that by comparing $t=1$ and $t=4$, one can see that the policy becomes conservative over time. At the time $4$, the action is hardly affected by the price and just tries to close the position as much as possible, most likely due to the eventual penalty.
@@ 5.3 Performance comparisons
We analyze the total costs’ distribution under the policy of each model based on 300,000 out-of-sample simulations. The results are shown in Table 1, where the value of the statistics of the total costs is shown. RL-mean achieves the lowest mean cost (-0.524), but its $\mathrm{ES}_{0.8}$ (0.253) and $\mathrm{ES}_{0.6}$ (0.086) values are not necessarily the lowest, reflecting a trade-off between optimizing for the mean and risk-sensitive measures. RL-$\mathrm{ES}_{0.8}$ achieves the lowest $\mathrm{ES}_{0.8}$ of the total costs (0.215), which aligns with its optimization objective. RL-$\mathrm{ES}_{0.6}$ shows the lowest $\mathrm{ES}_{0.6}$ of the total costs (0.079), reflecting its focus on mitigating risks in the worst 40% of outcomes. Our proposed method better controls the Expected Shortfall value compared to the other two methods. Under different $\alpha$’s, the learned policies achieve the lowest $\mathrm{ES}_{\alpha}$ values, respectively. The method we propose has a clear advantage in addressing (16) and serves as a valuable tool for controlling tail risk.

### v2

@@ Abstract
Keywords: Risk-sensitive reinforcement learning, Actor-Critic algorithm, convex scoring function, augmented state process, two-stage optimization, error analysis
@@ 1 Introduction
However, due to the time-non-separable nature in (2), it is well known that the time-inconsistency issue arises so that the standard dynamic programming principle cannot be employed directly in devising the algorithm. Our technical and practical contributions are summarized below.
@@
Secondly, we develop an Actor-Critic (AC) algorithm to optimize the policy and value function using neural networks. The Critic-step ensures that the value function approximates the true value uniformly, while the Actor-step updates the policy based on the policy gradient. Theoretical results demonstrate the approximation of the algorithm and provide bounds on the approximation error for both the value function (Theorem 2) and the optimal auxiliary variable $\upsilon$ (Theorem 3).
@@ 2.2 Risk Measures and Scoring Functions
In the context of quantitative risk management, a scoring function is a metric to evaluate and compare competing risk forecasts. As a definition (Chapter 9 of McNeil et al., (2015)), a classical scoring function $f:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ satisfies:
  • for any $y$ and $v$, $f(y,\upsilon)\geqslant 0$ and $f(y,\upsilon)=0$ if and only if $y=\upsilon;$
  • for a fixed $y$, $f(y,\upsilon)$ is increasing in $\upsilon$ for $\upsilon>y$ and decreasing for $\upsilon<y;$
  • for a fixed $y$, $f(y,\upsilon)$ is continuous in $\upsilon.$
@@
Definition 1 is quite general because the convexity is closely tied to unimodality. If the convexity is replaced by the condition that $-f(y,\upsilon)$ is unimodal in $\upsilon$ (i.e., for a fixed $y$, there exists some $\upsilon_{0}\in\mathbb{R}$ such that $f(y,\cdot)$ is decreasing on $(-\infty,\upsilon_{0}]$ and increasing on $[\upsilon_{0},\infty)$), then the scoring function satisfies the second condition in the definition of the classical scoring function above; see also Bellini and Bignozzi, (2015). In our paper, we focus on the convexity-based formulation to enhance the clarity and facilitate some mathematical arguments. The next result for stochastic optimization readily follows from Definition 1.
@@ 3 Problem Formulation
where the function $f$ is a convex scoring function and $h$ is a bivariate transform function strictly increasing w.r.t. its first argument, and the second argument $\upsilon$ is the auxiliary variable. Such risk measures are not required to be coherent or even monetary, but they encompass a broad class of functionals and many cases of coherent risk measures, such as expectation $\rho(Y)=\inf_{\upsilon\in\mathbb{R}}\mathbb{E}[Y]$ and Expected Shortfall (ES) $\rho(Y)=\inf_{\upsilon\in\mathbb{R}}\mathbb{E}[\upsilon+\frac{1}{\alpha}(Y-\upsilon)^{+}].$ Many convex (but may not be coherent) risk measures are also covered, such as entropic risk measures $\rho(Y)=\inf_{\upsilon\in\mathbb{R}}\frac{1}{\gamma}\log\left(\mathbb{E}[\exp(\gamma Y)]\right)$ with a risk-aversion parameter $\gamma>0$ and expected utility $\rho(Y)=\inf_{\upsilon\in\mathbb{R}}\mathbb{E}[u(Y)]$ with a general convex function $u$ (see Rockafellar and Uryasev, (2013)).
Note that some risk measures in our consideration are not even monetary, such as median absolute deviation (MAD) $\rho(Y)=\inf_{\upsilon\in\mathbb{R}}\mathbb{E}[|Y-\upsilon|],$ variance $\rho(Y)=\inf_{\upsilon\in\mathbb{R}}\mathbb{E}[|Y-\upsilon|^{2}]$ and asymmetric variance $\rho(Y)=\inf_{\upsilon\in\mathbb{R}}\,\mathbb{E}\left[\alpha(Y-\upsilon)^{2}\cdot\mathds{1}_{\left\{Y\geqslant\upsilon\right\}}+(1-\alpha)(Y-\upsilon)^{2}\cdot\mathds{1}_{\left\{Y\leqslant\upsilon\right\}}\right]$. Another important class of risk measures is the family of mean-risk utilities in Example 1, which are constructed as the combination of an expectation and a risk penalty.
@@ 4 Methodology
We need to further restrict the range of $\upsilon$ to a compact set such that it contains at least one optimal solution $\upsilon^{*}$ of Problem (7). To this end, we first show in Proposition 1 below that the range of $\upsilon$ can be restricted to a compact set. However, in some cases, the conditions in Proposition 1 may not be satisfied (see Remark 3) and we have to assume that the minimum of the objective is attained on $[\underline{\upsilon},\bar{\upsilon}]$ for some pre-known $\underline{\upsilon},\bar{\upsilon}\in\mathbb{R}.$ This assumption is commonly used in some existing studies in RL; see Baxter and Bartlett, (2001). Note that even when this assumption is not fulfilled, we can still find $\varepsilon$-correct range for $\upsilon,$ i.e., that for any $\varepsilon>0,$ there exists an interval $[\underline{\upsilon}^{\varepsilon},\bar{\upsilon}^{\varepsilon}]$ such that
@@
for any random variable $Y$ satisfying $\underline{y}\leqslant Y\leqslant\bar{Y}$ almost surely. To address a similar issue, Ni and Lai, (2022) derives a lower bound of the optimal $\upsilon$ (which agrees with Remark 3) and imposes an assumption on its upper bound for the entropic VaR.
@@
Consider the entropic VaR mentioned in Example 2. With $f(y,\upsilon)=\frac{1}{\alpha}\exp(e^{\upsilon}y)$, $h(x,\upsilon)=e^{-\upsilon}\log(x)$ and $g(x)=\exp(x),$ condition (1) holds for the entropic VaR if $[\underline{y},\bar{y}]\subset[0,\infty)$. However, conditions (2) and (3) are not satisfied because $g(h(f(y,\upsilon),\upsilon))$ is strictly decreasing for $\upsilon\in\mathbb{R}$ for fixed $y.$ In this case, for any $\varepsilon>0,$ we can find $[\underline{\upsilon}^{\varepsilon},\bar{\upsilon}^{\varepsilon}]$ such that
@@
The idea behind this formula is similar to the approach proposed by Bäuerle and Ott, (2011) to reduce the MDP under ES. However, their approach relies on the specific form of $(y-\upsilon)^{+}$ in the scoring function for ES, which facilitates the simple fusion of $Y$ and $\upsilon$ into one state process. For the convex scoring function, their approach no longer holds. However, our proposed method separates the variable $\upsilon$ and the augmented state process $Y$ such that we can devise the RL algorithm to accommodate more types of risk measures.
@@
Combining Theorem 1, the dominated convergence theorem and the continuity of $h,$ the next proposition holds.
@@
Therefore, studying the properties of $\mathbb{E}[V^{*}_{0,\upsilon}(S_{0},0)]$ is of fundamental importance. Proposition 3 establishes the continuity of $\upsilon\mapsto\mathbb{E}[V^{*}_{0,\upsilon}(S_{0},0)]$, a key result in showing the approximation of our RL algorithm in later subsections. Notably, our analysis does not require a continuity assumption on the MDP, rendering the proof of Proposition 3 technically nonstandard.
@@
We have that $\upsilon\mapsto\mathbb{E}[V^{*}_{0,\upsilon}(S_{0},0)]$ is continuous on $\Upsilon$ and $\upsilon\mapsto h(\mathbb{E}[V^{*}_{0,\upsilon}(S_{0},0)],\upsilon)$ has a minimizer (not necessarily unique) on $\mathbb{R}.$ Furthermore,
@@ 4.1 Part 1: Gradient Calculation and Neural Network (NN) Training
We adopt the policy gradient Actor-Critic RL algorithm for the inner optimization problem; see Konda and Tsitsiklis, (2003). That is, for the inner dynamic optimization problem, we alternatively learn two functions, the value function $V$ and the policy $\pi.$ In our work, both the value function and the policy are characterized by fully connected and multilayered feed-forward networks, for which the parameters are denoted by $\theta\in\Theta$ and $\phi\in\Phi$, respectively. The inputs to both networks include $(t,\upsilon,s,y).$ The network’s outputs are continuous with respect to $(t,\upsilon,s,y)$; see also some existing studies using the policy gradient approach on SRP in Tamar et al., (2016), Zhang et al., (2021), Ni and Lai, (2022) and Coache and Jaimungal, (2024). In contrast to these studies, our algorithm does not require the simulation-upon-simulation approach or the sample-average-approximation approach, which greatly reduces the computational complexity in simulating trajectories.
@@ 4.3 Part 3: Sampling the auxiliary variable $\upsilon$
Perform L(n) Critic-steps and get $\text{Perform $L(n)$ Critic-steps and get $\phi_{n}$};$
Perform L(n) Actor-steps and get $\text{Perform $L(n)$ Actor-steps and get $\theta_{n}$};$
@@
we have $\lim_{n\to\infty}h(\mathbb{E}[V^{\pi_{n,\varepsilon}}_{0,\upsilon_{n+1}}(S_{0},0)],\upsilon_{n+1})$ exists finitely. Furthermore, if there exists a nonnegative function $\delta:\Upsilon\times\Upsilon\to[0,\infty)$ such that for any $\upsilon^{\prime}\in\Upsilon,\ \pi\in\tilde{\Pi},$ the following two conditions hold:
@@
Algorithm 4 Complete Recipe based on a Convex Scoring Function
@@
4.3.1 Expected Shortfall (ES) Case
$\mathrm{ES}$ is widely used in risk-sensitive RL (see Bäuerle and Ott, (2011), Chow et al., (2015), Tamar et al., (2015), Kashima, (2007), Du et al., (2022), Godbout et al., (2021)). Our approach can handle the $\mathrm{ES}$ objective by taking $f(y,\upsilon)=\upsilon+(y-\upsilon)^{+}$ and $h(x,\upsilon)=x.$ We present here that under the $\mathrm{ES}$ objective, the sampling method described above can be further simplified to avoid the iterative use of SGD to solve (14). For any initial $\upsilon^{*}_{0}$ and parameters $\phi_{0}$ and $\theta_{0},$ we sample $\upsilon$ and update the parameters according to the following process: in stage $n\ (n\geqslant 1),$ we
@@ 4.3.1 Expected Shortfall (ES) Case
Perform L(n) Critic-steps and get $\text{Perform $L(n)$ Critic-steps and get $\phi_{n}$};$
Perform L(n) Actor-steps and get $\text{Perform $L(n)$ Actor-steps and get $\theta_{n}$};$
@@
It holds that $\lim_{n\to\infty}\mathbb{E}[V^{\pi_{n,\varepsilon}}_{0,\upsilon_{n+1}}(S_{0},0)]$ exists finitely. Furthermore, if there exists some $\Delta>0$ such that for any $\upsilon^{\prime}\in\Upsilon$ and $\pi\in\tilde{\Pi},$ the following two conditions hold:
@@ 5 Application: Statistical Arbitrage
  • RL-$\mathrm{Mean}$-$\mathrm{Var}:$ our proposed RL model minimizing the mean-variance utility, i.e. $\rho(Y)=\mathbb{E}[Y]+\lambda\mathbb{E}[(Y-\mathbb{E}[Y])^{2}],$ where $\lambda$ is chosen as $1$;
@@ 5.2 Policy Interpretation (Initial and Dynamic)
Figure 4 shows the policy learned after time $0$ Since the proposed method treats $y$ as a new (augmented) state, to understand the learned strategy, we propose a three-dimensional heat map illustration, which represents the action on different $(y,S,q)$ triples. Similar to time $0,$ the agent still suggests buying when the price is low and selling when the price is high again, but the policy becomes conservative as the value of $y$ increases. This can be illustrated with an example at $t=2$ in Figure 4, noting that when the price is very low (close to 0.5), with the same inventory $Q=2,$ the agent buys when the $y$ value is smaller and tends to maintain the position when the $y$ value is larger. Another point is that by comparing $t=1$ and $t=4$, one can see that the policy becomes conservative over time. At the time $4$, the action is hardly affected by the price and just tries to close the position as much as possible, most likely due to the terminal-time penalty.
@@ 5.3 Performance comparisons
We analyze the total costs’ distribution under the policy of each model based on 300,000 out-of-sample simulations. The results are shown in Table 1, where the value of the statistics of the total costs is shown. RL-mean achieves the lowest mean cost (-0.524), but its $\mathrm{ES}_{0.8}$ (0.253) and $\mathrm{ES}_{0.6}$ (0.086) values are not necessarily the lowest, reflecting a trade-off between optimizing for the mean and risk measures. RL-$\mathrm{ES}_{0.8}$ achieves the lowest $\mathrm{ES}_{0.8}$ of the total costs (0.215), which aligns with its optimization objective. RL-$\mathrm{ES}_{0.6}$ shows the lowest $\mathrm{ES}_{0.6}$ of the total costs (0.079), reflecting its focus on mitigating risks in the worst 40% of outcomes. Our proposed method better controls the Expected Shortfall value compared to the other two methods. Under different $\alpha$’s, the learned policies achieve the lowest $\mathrm{ES}_{\alpha}$ values, respectively. The method we propose has a clear advantage in addressing (16) and serves as a valuable tool for controlling tail risk.